In an isosceles triangle, there are two sides, called legs, with the same length. The third side
is called the base. If an isosceles triangle has perimeter 345 cm and base lenght 85 cm, what
is the length of each leg?

Answer by oberobic(2304)

L = Length of each leg
B = Length of base = 85
P = 2L + B = 345
2L + 85 = 345
Subtract 85 from both sides
2L = 345 - 85 = 260
Divide by 2
L = 130